Bluestem Supply does not does not segregate sales and sales taxes at the time of sale. The register total for March 16 is $10,388. All sales are subject to a 6% sales tax.


Compute sales taxes payable.

Sales tax payable $Bluestem Supply does not does not segregate sales


Make the entry to record sales taxes payable and sales. (Credit account titles are automatically indented when amount is entered. Do not indent manually.)


Date
Account Titles and Explanation
Debit
Credit
Mar. 16
Bluestem Supply does not does not segregate sales
Bluestem Supply does not does not segregate sales
Bluestem Supply does not does not segregate sales
Bluestem Supply does not does not segregate sales
Bluestem Supply does not does not segregate sales
Bluestem Supply does not does not segregate sales
Bluestem Supply does not does not segregate sales
Bluestem Supply does not does not segregate sales
Bluestem Supply does not does not segregate sales
Answer #1
Concepts and reason

Sales tax: Sales tax or value added tax is a type of tax which is levied on goods at the point of sale. Buyer of the goods must bear the burden of the sales tax on the goods purchased by him. Duty to collect, compliance with the provision of sales tax and deposit of sales tax is on the seller of the goods. Rate of sales tax may be different in different jurisdictions.

Sales tax payable: Duty to collect the sales tax is on the seller, therefore seller at the time of the sale collects the sales tax from the buyer along with the consideration of sales. Until the sales taxes are deposited with the governing body, it is recorded as the sales tax liability.

Fundamentals

Sales: Sales is the amount of revenue earned by a particular organization through primary activities. It is the amount presented in the income summary account. Expenses are reduced from the sales amount to compute the net income or loss for a particular period.

Journal entries: Journal entry records the accounting transactions of a business in a journal book. All the business transactions are recorded in the chronological order using the double entry system of accounting.

Calculate the amount of sales tax payable as shown below:

Salestaxpayable=Totalamount×Salestaxrate100+Salestaxrate=$10,388×6%100%+6%=$10,388×6106=$588\begin{array}{c}\\{\rm{Sales tax payable}} = {\rm{Total amount}} \times \frac{{{\rm{Sales tax rate}}}}{{100 + {\rm{Sales tax rate}}}}\\\\ = \$ 10,388 \times \frac{{6\% }}{{100\% + 6\% }}\\\\ = \$ 10,388 \times \frac{6}{{106}}\\\\ = \$ 588\\\end{array}

Post Ref.
Date
Mar, 16
DR.(S)
10.388
CR.(S)
Accounts
Cash
Sales tax payable
Sales
(To record the sales tax payable and sales)

Working note:

Calculate amount of sales exclusive of sales tax:

Amountofsalesexclusiveofsalestax=SalesinclusiveofsalestaxAmountofsalestax=$10,388$588=$9,800\begin{array}{c}\\{\rm{Amount of sales exclusive of sales tax = Sales inclusive of sales tax - Amount of sales tax}}\\\\{\rm{ = \$ 10,388 - \$ 588}}\\\\{\rm{ = \$ 9,800}}\\\end{array}

Ans:

Sales tax payable is $588.

Post Ref.
Date
Mar, 16
DR.(S)
10.388
CR.(S)
Accounts
Cash
Sales tax payable
Sales
(To record the sales tax payable and sales)